# DZ-TDPO: Non-Destructive Temporal Alignment for Mutable State Tracking in Long-Context Dialogue

### v1

@@ Abstract
Long-context dialogue systems suffer from State Inertia, where static constraints prevent models from resolving conflicts between evolving user intents and established historical context. To address this, we propose DZ-TDPO, a non-destructive alignment framework that synergizes conflict-aware dynamic KL constraints with a learnable temporal attention bias. Experiments on the Multi-Session Chat (MSC) dataset demonstrate that DZ-TDPO achieves state-of-the-art win rates (86.2% on Phi-3.5) while maintaining robust zero-shot generalization. Crucially, our scaling analysis reveals a "Capacity-Stability Trade-off": while smaller models incur an "alignment tax" (perplexity surge) to overcome historical inertia, the larger Qwen2.5-7B model achieves near-perfect alignment (99.4% win rate) with negligible perplexity overhead. This confirms that TAI can be alleviated via precise attention regulation rather than destructive weight updates, preserving general capabilities (MMLU) across model scales. Code and data are available: https://github.com/lyj20071013/DZ-TDPO
@@ 1 Introduction
Large Language Models (LLMs) have witnessed a rapid expansion in tackling long-sequence problems, driven by efficient fine-tuning techniques like LongLORA (Chen et al., 2024a) and positional interpolation methods such as YaRN (Peng et al., 2023). Existing architectures have achieved remarkable success in "Static Retrieval" tasks, where the goal is to locate a specific piece of information within a vast context window (e.g., "Needle-in-a-Haystack" tests). In these scenarios, information is additive and non-conflicting, and the primary challenge is effectively extending the receptive field to "remember" potentially infinite history.
@@
Despite the success of Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022), standard alignment methods like Direct Preference Optimization (DPO) (Rafailov et al., 2023) struggle to resolve these dynamic conflicts. We argue that standard DPO imposes a "Static Alignment Constraint" that treats all historical tokens as immutable priors. Consequently, when a model attempts to update its state to match a recent turn, it incurs a heavy penalty for deviating from the reference model’s historical behavior. Correcting this inertia often requires aggressive parameter updates, leading to a significant "Alignment Tax" Amanda Askell (2021); Lin et al. (2024)—a degradation in general linguistic capabilities, manifested as a catastrophic surge in perplexity (PPL) and a loss of structural coherence.
@@
  • Scaling Insight: We provide the first analysis of the Capacity-Stability Trade-off in temporal alignment. Experiments on Qwen2.5-7B Yang et al. (2024) show that larger models can internalize temporal bias with minimal "Alignment Tax" , contrasting with the steeper cost paid by smaller models, thus offering a scalable solution for long-context agents.
@@ 2.1 Related Work
To handle long-sequence inputs, the community has proposed various structural innovations, such as RoPE Su et al. (2024) for position encoding and ALiBi Press et al. (2022) for length extrapolation, enabling models to process sequences exceeding 100k tokens. Beyond architecture, Approaches like LongAlign Bai et al. (2024) and LongPO Guanzheng Chen (2025) focus on extending the receptive field to address retrieval tasks ("Finding the needle"). In contrast, DZ-TDPO addresses the conflict task ("Updating the needle"), solving the State Inertia that these capacity-focused methods neglect, often operating under the assumption that all historical context is potentially relevant. Recent efficiency-focused works have explored sparse attention mechanisms: StreamingLLM (Xiao et al., 2024) identifies “attention sinks” to maintain generation stability, while H2O (Zhang et al., 2023) evicts non-heavy-hitter tokens to reduce KV cache footprint. Similarly, Ring Attention (Liu et al., 2023) optimizes computation for near-infinite contexts. However, these methods primarily focus on computational efficiency or retrieval recall. In contrast, DZ-TDPO addresses the conflict resolution problem. Recent studies have identified distinct attention phenomena, such as ‘attention sinks’ Xiao et al. (2024) that disproportionately weight initial tokens. Furthermore, benchmarks like RULER Hsieh et al. (2024) highlight that effectively utilizing long context for precise state tracking remains a significant challenge. In contrast, DZ-TDPO focuses on the conflict resolution problem (“which part should it trust?”). Unlike general long-context methods, our approach specifically addresses the decision dilemma when historical information contradicts the current state—a dimension rarely discussed in prior literature.
@@ 3.1 Optimization Level: TDPO-DKL
Unlike standard exponential decay, we argue that the decay rate should depend on the semantic conflict between the current user input and history. We map dialogue turns into a latent semantic space and define the adaptive decay temperature $\tau(u_{T})$ for the current user turn as:
@@
where $e_{T}$ and $e_{i}$ are sentence embeddings encoded by a lightweight Transformer (SBERT). A high cosine similarity implies a potential state update or topic revisitation, triggering a lower $\tau$ to sharpen the model’s focus on the present.
@@
$\beta(t;T)=\beta_{0}\cdot\left[\alpha+(1-\alpha)\cdot\exp\left(-\frac{T-t}{\tau(x_{T})}\right)\right]$ (5)
@@
$w(t;T)=\exp\left(-\frac{T-t}{\tau(x_{T})}\right)$ (6)
@@
We first define the implicit log-ratio margin $\mathcal{M}_{\theta}(x_{t},y_{w},y_{l})$ as:
$\mathcal{M}_{\theta}(x_{t},y_{w},y_{l})=\log\frac{\pi_{\theta}(y_{w}|x_{t})}{\pi_{\text{ref}}(y_{w}|x_{t})}-\log\frac{\pi_{\theta}(y_{l}|x_{t})}{\pi_{\text{ref}}(y_{l}|x_{t})}$ (7)
@@
$\mathcal{L}_{\text{TDPO-DKL}}(\theta)=-\mathbb{E}_{\mathcal{D}}\left[w(t;T)\cdot\log\sigma\left(\beta(t;T)\cdot\mathcal{M}_{\theta}(x_{t},y_{w},y_{l})\right)\right]$
@@ 3.2 Representation Level: Dual-Zone Temporal Attention
Building on the robust DZ-TA, we structurally enforce the distinction between core values and evolving history. We conceptualize the context window $C$ as consisting of two distinct regions:
We conceptualize the context window $C$ as consisting of two distinct regions: the Immutable Anchor Zone ($Z_{anchor}$) (indices 0 to $k$), covering the System Prompt and core safety guidelines which must remain invariant; and the Mutable State Zone ($Z_{state}$) (indices $k+1$ to $T$), covering the conversational history subject to plasticity.
@@
Here, $\lambda$ is the shared learnable scalar parameter initialized to 0.5, which autonomously optimizes the "forgetting rate" for user history.
@@ 4.1 Experimental Setup
Implementation Details
For TDPO-DKL, we set the base KL coefficient $\beta_{0}=0.1$ and the minimum constraint ratio $\alpha=0.3$. To enable the Conflict-Aware Adaptive Decay, we set the base temporal horizon $\tau_{base}=8.0$, the scaling factor $\gamma=0.8$, and the minimum decay floor $\tau_{min}=0.5$. For the adaptive decay mechanism, we utilize the all-MiniLM-L6-v2 model, a distilled Transformer based on the MiniLM architecture (Wang et al., 2020). We execute semantic encoding using the Sentence-BERT framework (Reimers and Gurevych, 2019) to compute the cosine similarity between the current turn and historical context. This approach overcomes the limitations of surface-level lexical overlap (as discussed in Limitations) by capturing latent semantic contradictions (e.g., "Vegan" vs "Steak"). Training is conducted on a single NVIDIA A800 GPU with a batch size of 16 for 3 epoch to prevent overfitting.
@@ 4.2 Results
Base Model 20.2 % 10.6 % 22.1 5.27
Standard DPO 52.2 % 63.2 % 124.1 5.35
SimPO 60.8 % 30.8  % 99.6 5.28
TDPO-DKL (w/o DZ-TA) 76.4 % 68.1 % 100.9 5.67
DZ-TDPO (Ours) 86.2 % 71.0 % 24.8 5.45
Table 1: For Standard DPO, we report the performance at Epoch 1, as we observed severe reward hacking and perplexity degradation (>100) in subsequent epochs. For SimPO, we report the performance at Epoch 2.
Evaluation of Mutable State Tracking. As shown in Table 1, purely optimization-based baselines struggle profoundly with the State Inertia inherent in long-context dialogues.
A critical observation is the catastrophic behavior of Standard DPO. While it achieves a baseline win rate of 52.2%, it incurs a massive "Alignment Tax," with validation perplexity exploding to 124.1. We argue that this is not a training artifact, but a theoretical inevitability of the Static Alignment Constraint. When the user updates their state (e.g., $A\rightarrow\neg A$), standard DPO drives the model to maximize the likelihood of $\neg A$ (the new truth), yet the static KL constraint anchors it to $A$ (the historical prior). Without an attention mechanism to resolve this contradiction, the model is forced to shatter its pre-trained linguistic priors to satisfy the reward objective, resulting in distribution collapse.
In stark contrast, DZ-TDPO achieves state-of-the-art performance with an 86.2% in-domain win rate. Most importantly, it maintains a healthy perplexity of 24.8, which is comparable to the Base Model (22.1) and significantly lower than the optimization-based baselines. This confirms the efficacy of our Dual-Zone Architecture: by structurally suppressing the attention mass of outdated states via the Mutable State Zone ($Z_{state}$), the optimization module (TDPO-DKL) operates on a "clean" gradient landscape. The model learns to override the state without destroying its general linguistic capabilities.
SimPO, while avoiding reference model constraints, suffers from severe overfitting, with OOD performance on UltraChat plummeting to 30.8%. Conversely, DZ-TDPO achieves a robust 71.0% OOD win rate. This validates that our DZ-TA acts as a generalized low-rank regularizer, learning a universal "recency policy" rather than memorizing specific dataset patterns.
Beyond preference win rates, we evaluate the generative quality of the models using standard n-gram metrics including SacreBLEU (Post, 2018), ROUGE (Lin, 2004) and BERTScore (Zhang et al., 2020) against the reference responses. (detailed n-gram metrics provided in Appendix C.2)
@@ 4.3 The Capacity-Stability Trade-off
To investigate the impact of model scale, we extended our experiments to the Qwen2.5-7B-Instruct model. As shown in Table 2, DZ-TDPO demonstrates consistent scalability, though a "Capacity-Efficiency" pattern emerges.
The 7B model achieves a near-perfect Win Rate of 99.4% on the MSC dataset, surpassing the 3.8B model’s 86.2%. Crucially, the perplexity analysis reveals the "cost" of this alignment. While the smaller Phi-3.5 (3.8B) incurs a slight "stability cost" ($\Delta$ PPL +2.7) to accommodate the temporal bias, the 7B model absorbs this mechanism with negligible overhead ($\Delta$ PPL +1.95). This suggests that while DZ-TDPO is non-destructive for both, larger models possess sufficient parametric redundancy to internalize temporal dynamics more efficiently.
Beyond in-domain performance, the 7B model exhibits superior zero-shot generalization. On UltraChat (OOD), it attains a 91.8% win rate at the standard 4k context. Furthermore, we evaluated length extrapolation by extending the context to 8k four times the training window. The model retains a robust 78.0% win rate, confirming that the DZ-TA mechanism learns a generalized distance penalty rather than overfitting to specific positions, avoiding catastrophic "contextual myopia."
Both models maintain general world knowledge, with minimal perplexity variation on MMLU ($\Delta$ +0.18 for 3.8B vs +0.46 for 7B). This confirms that DZ-TDPO resolves temporal conflicts via precise attention regulation, serving as a safe alignment solution across model scales.
@@
Win Rate $\uparrow$ 86.2% 99.4%
Alignment Tax (PPL $\Delta$) $\downarrow$ +2.7 +1.95
@@
Win Rate (4k Context) $\uparrow$ 71.0% 91.8%
Win Rate (8k Extrapolation) $\uparrow$ 48.4% 78.0%
@@ 4.4 Qualitative Analysis
To rigorously rule out "Contextual Myopia" (where $\lambda$ might aggressively suppress history), we conducted a Non-Conflicting Needle-in-a-Haystack evaluation. As shown in Table 5 (Appendix B.4), despite being trained on contexts <2.4k<2.4k, the model successfully retrieved specific entities across 2k–64k intervals. This validates that $\lambda\approx 0.68$ acts as a soft semantic filter, suppressing background noise while propagating strong, non-conflicting signals to enable robust extrapolation.
@@ 5 Discussion
The Theoretical Inevitability of the "Alignment Tax"Our findings reveal a fundamental mechanism mismatch in standard alignment algorithms when applied to mutable state tracking. Standard DPO, enforced by a static KL constraint, inherently treats all historical tokens as immutable priors. The catastrophic perplexity surge observed in the Standard DPO baseline (124.1) is not a training artifact, but empirical evidence of State Inertia. When a user updates their state (e.g., $A\rightarrow\neg A$), the static constraint forces the model to maximize the likelihood of $\neg A$ while the entire history implies $A$. Without an attention mechanism to resolve this contradiction, the model is compelled to "shatter" its pre-trained linguistic manifolds to satisfy the conflicting reward, resulting in distribution collapse. DZ-TDPO resolves this by decoupling the alignment process. The Dual-Zone Architecture acts as a representation-level filter, proactively suppressing the signal of outdated states. Consequently, the optimization module (TDPO-DKL) operates on a "clean" gradient landscape, allowing the model to internalize the state update without paying the "Alignment Tax." This explains why our method achieves SOTA win rates while maintaining near-baseline perplexity.

### v2

@@ Abstract
Long-context dialogue systems suffer from State Inertia, where static constraints prevent models from resolving conflicts between evolving user intents and established historical context. To address this, we propose DZ-TDPO, a non-destructive alignment framework that synergizes conflict-aware dynamic KL constraints with a calibrated temporal attention bias. Experiments on the Multi-Session Chat (MSC) dataset demonstrate that DZ-TDPO achieves state-of-the-art win rates (55.4% on Phi-3.5) while maintaining robust zero-shot generalization. Our scaling analysis reveals a "Capacity-Stability Trade-off": while smaller models incur an "alignment tax" (perplexity surge) to overcome historical inertia, the larger Qwen2.5-7B model achieves 50.8% win rate with negligible perplexity overhead. This confirms that TAI can be alleviated via precise attention regulation rather than destructive weight updates, preserving general capabilities (MMLU) across model scales. Code and data are available: https://github.com/lyj20071013/DZ-TDPO
@@ 1 Introduction
Large Language Models (LLMs) have rapidly advanced in tackling long-sequence problems, driven by efficient fine-tuning techniques like LongLORA (Chen et al., 2024a) and positional interpolation methods such as YaRN (Peng et al., 2023). Existing architectures have achieved remarkable success in "Static Retrieval" tasks, where the goal is to locate a specific piece of information within a vast context window (e.g., "Needle-in-a-Haystack" tests). In these scenarios, information is additive and non-conflicting, and the primary challenge is effectively extending the receptive field to "remember" potentially infinite history.
@@
Despite the success of Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022), standard alignment methods like Direct Preference Optimization (DPO) (Rafailov et al., 2023) struggle to resolve these dynamic conflicts. We argue that standard DPO imposes a "Static Alignment Constraint" that treats all historical tokens as immutable priors. Consequently, when a model attempts to update its state to match a recent turn, it incurs a heavy penalty for deviating from the reference model’s historical behavior. Correcting this inertia often requires aggressive parameter updates, leading to a significant "Alignment Tax" Askell et al. (2021); Lin et al. (2024)—a degradation in general linguistic capabilities, manifested as a catastrophic surge in perplexity (PPL) and a loss of structural coherence.
@@
  • Scaling Insight: We provide the first analysis of the Capacity-Stability Trade-off in temporal alignment. Experiments on Qwen2.5-7B Team et al. (2025) show that larger models can internalize temporal bias with minimal "Alignment Tax" , contrasting with the steeper cost paid by smaller models, thus offering a scalable solution for long-context agents.
@@ 2.1 Related Work
To handle long-sequence inputs, the community has proposed various structural innovations, such as RoPE Su et al. (2024) for position encoding and ALiBi Press et al. (2022) for length extrapolation, enabling models to process sequences exceeding 100k tokens. Beyond architecture, approaches like LongAlign Bai et al. (2024) and LongPO Chen et al. (2025) focus on extending the receptive field to address retrieval tasks ("Finding the needle"). In contrast, DZ-TDPO addresses the conflict task ("Updating the needle"), solving the State Inertia that these capacity-focused methods neglect, often operating under the assumption that all historical context is potentially relevant. Recent efficiency-focused works have explored sparse attention mechanisms: StreamingLLM (Xiao et al., 2024) identifies “attention sinks” to maintain generation stability, while H2O (Zhang et al., 2023) evicts non-heavy-hitter tokens to reduce KV cache footprint. Similarly, Ring Attention (Liu et al., 2023) optimizes computation for near-infinite contexts. However, these methods primarily focus on computational efficiency or retrieval recall. In contrast, DZ-TDPO addresses the conflict resolution problem. Recent studies have identified distinct attention phenomena, such as ‘attention sinks’ Xiao et al. (2024) that disproportionately weight initial tokens. Furthermore, benchmarks like RULER Hsieh et al. (2024) highlight that effectively utilizing long context for precise state tracking remains a significant challenge. In contrast, DZ-TDPO focuses on the conflict resolution problem (“which part should it trust?”). Unlike general long-context methods, our approach specifically addresses the decision dilemma when historical information contradicts the current state—a dimension rarely discussed in prior literature.
@@ 3.1 Optimization Level: TDPO-DKL
Unlike in standard exponential decay methods, we argue that the decay rate should depend on the semantic conflict between the current user input and history. We map dialogue turns into a latent semantic space and define the adaptive decay temperature $\tau(u_{T})$ for the current user turn as:
@@
where $\mathbf{e}_{T}$ and $\mathbf{e}_{i}$ are sentence embeddings encoded by a lightweight Transformer (SBERT). A high cosine similarity implies a potential state update or topic revisitation, triggering a lower $\tau$ to sharpen the model’s focus on the present. Conversely, low similarity implies orthogonal topics where history should be preserved.
@@
$\beta(t;T)=\beta_{0}\cdot\left[1-(1-\alpha)\cdot\exp\left(-\frac{T-t}{\tau(u_{T})}\right)\right]$ (5)
@@
$w(t;T)=\exp\left(-\frac{T-t}{\tau(u_{T})}\right)$ (6)
@@
We first define the implicit log-ratio margin $\mathcal{M}_{\theta}(c,y_{w},y_{l})$ as:
$\mathcal{M}_{\theta}(c,y_{w},y_{l})=\log\frac{\pi_{\theta}(y_{w}|c)}{\pi_{\text{ref}}(y_{w}|c)}-\log\frac{\pi_{\theta}(y_{l}|c)}{\pi_{\text{ref}}(y_{l}|c)}$ (7)
@@
$\mathcal{L}_{\text{TDPO-DKL}}(\theta)=-\mathbb{E}_{\mathcal{D}}\left[w(t;T)\cdot\log\sigma\left(\beta(t;T)\cdot\mathcal{M}_{\theta}(c,y_{w},y_{l})\right)\right]$
@@ 3.2 Representation Level: Dual-Zone Temporal Attention
We conceptualize the context window $C$ as consisting of two distinct regions: the Immutable Anchor Zone ($Z_{anchor}$) (indices 0 to $L_{anc}$) covering the System Prompt and core safety guidelines which must remain invariant; and the Mutable State Zone ($Z_{state}$) (indices $L_{anc}+1$ to $T$) covering the conversational history subject to plasticity.
@@
Here, $\lambda$ is the shared scalar parameter. While theoretically learnable (see Appendix A.2), we fixed $\lambda=0.5$ as a structural prior (theoretical justification in Appendix A.2). This imposes a strong structural prior that enforces the ’forgetting rate’ for user history without introducing optimization instability.
@@ 4 Experiments
For preference evaluation, we employ DeepSeek-V3.2 (DeepSeek-AI et al., 2025) as the automated judge. Recent studies indicate that DeepSeek-V3.2 exhibits high correlation with human judgment and strong capabilities in complex reasoning tasks. We conduct pairwise comparisons where the judge is presented with the dialogue history, the specific instruction, and two anonymized model responses, following the standard LLM-as-a-Judge protocol (detailed protocol in Appendix G).
@@ 4.1 Experimental Setup
Detailed hyperparameters and training configurations are provided in Appendix B.2. Training is conducted on a single NVIDIA A800 GPU with a batch size of 32 for 6 epochs to prevent overfitting.
@@ 4.2 Results
Base Model - % - % 22.1 5.27
Standard DPO 45.8 % 49.2 % 102.3 5.28
SimPO 46.4 % 48.8  % 101.2 5.29
TDPO-DKL (w/o DZ-TA) 50.2 % 49.8 % 101.1 5.42
DZ-TDPO (Ours) 55.4 % 53.5 % 26.0 5.34
Table 1: For Standard DPO, we report the performance at Epoch 6, as we observed severe reward hacking and perplexity degradation (>100) in subsequent epochs. For SimPO, we report the performance at Epoch 24 (batch size of 128).
As shown in Table 1, purely optimization-based baselines struggle profoundly with the State Inertia inherent in long-context dialogues.
A critical observation is the catastrophic behavior of Standard DPO. While it achieves a baseline win rate of 45.8%, it incurs a massive "Alignment Tax," with validation perplexity exploding to 102.3. We argue that this is not a training artifact, but a theoretical inevitability of the Static Alignment Constraint. When the user updates their state (e.g., $A\rightarrow\neg A$), standard DPO drives the model to maximize the likelihood of $\neg A$ (the new truth), yet the static KL constraint anchors it to $A$ (the historical prior). Without an attention mechanism to resolve this contradiction, the model is forced to shatter its pre-trained linguistic priors to satisfy the reward objective, resulting in distribution collapse.
In stark contrast, DZ-TDPO achieves state-of-the-art performance with an 55.4% in-domain win rate. Most importantly, it maintains a healthy perplexity of 26.0, which is comparable to the Base Model (22.1) and significantly lower than the optimization-based baselines. This confirms the efficacy of our Dual-Zone Architecture: by structurally suppressing the attention mass of outdated states via the Mutable State Zone ($Z_{state}$), the optimization module (TDPO-DKL) operates on a "clean" gradient landscape. The model learns to override the state without destroying its general linguistic capabilities.
SimPO, while avoiding reference model constraints, demonstrates limited zero-shot generalization, with OOD performance on UltraChat recording 48.8%. Conversely, DZ-TDPO achieves a robust 53.5% OOD win rate. This suggests that our DZ-TA acts as a generalized low-rank regularizer, capturing a universal ‘recency principle’ rather than over-optimizing for in-domain patterns.
Beyond preference win rates, we evaluate the generative quality of the models using standard n-gram metrics including SacreBLEU (Post, 2018), ROUGE (Lin, 2004) and BERTScore (Zhang et al., 2020) against the reference responses. (detailed n-gram metrics provided in Appendix C.2). Furthermore, head-to-head comparisons (Appendix C.4) confirm DZ-TDPO’s superiority.
@@ 4.3 The Capacity-Stability Trade-off
To investigate the impact of model scale, we extended our experiments to the Qwen2.5-7B-Instruct model. As shown in Table 2, a distinct "Stability-Plasticity" pattern emerges compared to the smaller Phi-3.5 model.
Interestingly, the 7B model achieves a win rate of 50.8% on the MSC dataset, which is slightly lower than the 3.8B model’s 55.4%. We attribute this to the stronger Parametric Inertia inherent in larger models. With significantly more pre-trained knowledge, the 7B model possesses a more rigid "belief system," making it harder to override historical context with short-term updates compared to the more plastic 3.8B model.
However, the perplexity analysis reveals the advantage of this rigidity. While the smaller Phi-3.5 incurs a visible "stability cost" ($\Delta$ PPL +3.9+3.9) to accommodate the temporal bias, the 7B model absorbs the DZ-TDPO mechanism with negligible overhead ($\Delta$ PPL +1.95+1.95). This confirms that while larger models are harder to steer, they are also significantly more robust against distribution collapse. DZ-TDPO is proven to be non-destructive at scale, offering a safe alignment path.
Beyond in-domain performance, the 7B model maintains consistent zero-shot generalization. On UltraChat (OOD), it attains a 50.8% win rate, mirroring its in-domain behavior. Furthermore, general world knowledge remains intact, with minimal perplexity variation on MMLU ($\Delta$ +0.07 for 3.8B vs +0.45 for 7B). This scaling analysis suggests that resolving State Inertia in larger models may require stronger intervention strengths (e.g., lower $\tau$ or higher $\lambda$) to overcome their inherent parametric stability.
@@
Win Rate $\uparrow$ 55.4% 50.8%
Alignment Tax (PPL $\Delta$) $\downarrow$ +3.9 +1.95
@@
Win Rate (4k Context) $\uparrow$ 53.5% 50.83%
@@ 4.4 Qualitative Analysis
To rigorously rule out "Contextual Myopia" (where $\lambda$ might aggressively suppress history), we conducted a Non-Conflicting Needle-in-a-Haystack evaluation. As shown in Table 5 (Appendix B.5), despite being trained on contexts <2.4k<2.4k, the model successfully retrieved specific entities across 2k–8k intervals. This validates that $\lambda\approx 0.5$ acts as a soft semantic filter, suppressing background noise while propagating strong, non-conflicting signals to enable robust extrapolation.
@@ 5 Discussion
The Theoretical Inevitability of the "Alignment Tax"Our findings reveal a fundamental mechanism mismatch in standard alignment algorithms when applied to mutable state tracking. Standard DPO, enforced by a static KL constraint, inherently treats all historical tokens as immutable priors. The catastrophic perplexity surge observed in the Standard DPO baseline (102.3) is not a training artifact, but empirical evidence of State Inertia. When a user updates their state (e.g., $A\rightarrow\neg A$), the static constraint forces the model to maximize the likelihood of $\neg A$ while the entire history implies $A$. Without an attention mechanism to resolve this contradiction, the model is compelled to significantly disrupt its pre-trained linguistic features to satisfy the conflicting reward, resulting in distribution collapse. DZ-TDPO resolves this by decoupling the alignment process. The Dual-Zone Architecture acts as a representation-level filter, proactively suppressing the signal of outdated states. Consequently, the optimization module (TDPO-DKL) operates on a "clean" gradient landscape, allowing the model to internalize the state update without paying the "Alignment Tax." This explains why our method achieves SOTA win rates while maintaining near-baseline perplexity.